OpenStudy (anonymous):

the measure of the angles of a triangle are in the ratio 2:3:4. In degrees, the measure if the largest angle is ?

OpenStudy (michele_laino):

please, note that you have a chain of ratios. Now If I call with x, y and z the amplitudes of internal angles of your triangle, we can write: x+y+z=180 do you agree?

OpenStudy (michele_laino):

please note that I can not give you the solution directly, since the Code of Conduct

OpenStudy (michele_laino):

@YourFlyGuy

OpenStudy (michele_laino):

\[x+y+z=180°\] is it right?

Directrix (directrix):

>>angles of a triangle are in the ratio 2:3:4 Represent the angles as 2x, 3x, and 4x where x is a common factor of all 3 angles. 2x + 3x + 4x = 180 because the sum of the interior angles of a triangle is 180. 9x = 180 x = 20 The largest angle has measure 4*x. What will that be? @YourFlyGuy